A body travels 200 CM in the first 2 seconds and 220 CM in the next 5 second calculate the velocity at the end of the 7th second from the start

Grade:9

2 Answers

Arun
5 years ago

let initial velocity = u & accleration is a then

distance travelled in 5 sec = S7

using , s=ut + at2/2

          S7 = 7u + a49/2

               =  7u+24.5 a 

distance covered in 2 sec = S2 = 2u + 4a/2

                                           =2u+2a = 200           (given)

                                       u+a = 100             ........................1

now distance covered in 2 to 6 sec is given 220cm

      S7-200 = 220

     7u + 24.5a = 420       

        u + 3.5a = 60             .................2

from 1& 2

a = -16 & u=116

velocity after 7 sec = V7

       V7 = u + at

           =116 - 16*7=4cm/sec

Eshan
5 years ago
.